OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2.

OpenStudy (anonymous):

@jojo4eva please show your work

OpenStudy (anonymous):

Equation of the line with slope m and (x1,y1) as a point on it is given by y-y1=m(x-x1)

OpenStudy (anonymous):

yea so y-0=-2(x-3)

OpenStudy (anonymous):

yeah then y=-2x+6

OpenStudy (jojo4eva):

ya sorry i had it typed but my computer was giving me problems but @Princer_Jones is correct

OpenStudy (anonymous):

oh i see you multiplied 2 and 3

OpenStudy (anonymous):

a*(b+c)=a*b+a*c @omg1help